OpenStudy (akgorgeous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4

OpenStudy (sooobored):

both equations have different slopes so they cannot be parallel lines

OpenStudy (sooobored):

parallel lines will have the same slope

OpenStudy (mathmale):

You have only one "system of linear equations" here (not 2 or 3 or more). Your job is to determine whether or not you can solve this "system of linear equations." If your two equations represent 2 lines, and the lines have the same slope, then your lines never intersect, and thus you have no solution. "no x value" is incorrect; there'd be no y value either. Better to say, "no solution."

OpenStudy (mrnood):

the easiest way to visualise a straight line equation is to put it in slope intercept form y=mx+b your equations can be rewritten as y = 2/3 x-1 and y= 5/4 x -1 the value of m tells you the slope IF the 2 lines have different slope then they MUST intersect at one point only, and hence have one solution to the 'system'

OpenStudy (akgorgeous):

so x is -1

OpenStudy (mrnood):

how did you derive that? if oyu show some working you can be guided more easily

OpenStudy (mathmale):

How does it make sense to ask "so it's x=-1?" Again, two parallel lines do NOT intersect, and thus the system has NO solution. It is very important that you understand these principles.

OpenStudy (mrnood):

btw x=-1 is not correct

OpenStudy (akgorgeous):

What im trying to ask is if there is an x value and what it is. I thought the system was parallel therefore having no solution, x, or y value.

OpenStudy (mrnood):

IF the 2 lines have different slope then THEY ARE NOT PARALLEL and they MUST intersect at one point only, and hence have one solution to the 'system'

OpenStudy (mrnood):

You have confused the question by putting 'parallel' in the question please post it exactly as you have it

OpenStudy (akgorgeous):

that was the original question, i just updated it.

OpenStudy (mrnood):

y=mx+b the value of m tells you the slope your equations can be rewritten as y = 2/3 x-1 and y= 5/4 x -1 So WHAT is th e slope of each line - and are they equal?

OpenStudy (mrnood):

OK so now you are not starting with the assumption that they are parallel SO can you answer my question in my last post

OpenStudy (akgorgeous):

the slopes are different. so its not parallel

OpenStudy (mrnood):

correct now - the b value tells you where the lines cross the y axis where does each line cross?

OpenStudy (akgorgeous):

they both have the same y - intercept yea got it

OpenStudy (mrnood):

so where do they intersect each other?

OpenStudy (akgorgeous):

-1

OpenStudy (mrnood):

OK good but you need (x,y) point what is the x value at the y intercept?

OpenStudy (akgorgeous):

x=0 y=-1

OpenStudy (mrnood):

OK so you have your answer...